# Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey

### v1

@@ Abstract
Bipedal robots are garnering increasing global attention due to their potential applications and advancements in artificial intelligence, particularly in Deep Reinforcement Learning (DRL). While DRL has driven significant progress in bipedal locomotion, developing a comprehensive and unified framework capable of adeptly performing a wide range of tasks remains a challenge. This survey systematically categorizes, compares, and summarizes existing DRL frameworks for bipedal locomotion, organizing them into end-to-end and hierarchical control schemes. End-to-end frameworks are assessed based on their learning approaches, whereas hierarchical frameworks are dissected into layers that utilize either learning-based methods or traditional model-based approaches. This survey provides a detailed analysis of the composition, capabilities, strengths, and limitations of each framework type. Furthermore, we identify critical research gaps and propose future directions aimed at achieving a more integrated and efficient framework for bipedal locomotion, with potential broad applications in everyday life.
@@ I Introduction
Humans navigate complex and varied environments, performing diverse locomotion tasks with only two legs. To facilitate dynamic bipedal locomotion, model-based methods were introduced in the 1980s and have since evolved significantly [1, 2, 3]. These methods, characterized by rapid convergence, furnish a predictive framework for understanding environmental structures. However, they struggle to adapt in dynamically challenging environments that are difficult to model precisely. More recently, advancements in machine learning have provided novel approaches. Reinforcement learning (RL)-based methods, in particular, are adept at navigating the full dynamics of robot-environment interactions [4]. Additionally, hybrid approaches that combine model-based and learning-based methods have been developed to leverage the advantages of both. Yet, the question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
@@
To address this, we explore recent advancements in deep reinforcement learning (DRL)-based frameworks, categorizing control schemes into two primary types: (i) end-to-end and (ii) hierarchical. End-to-end frameworks map robot states directly to control outputs at the joint level, while hierarchical frameworks adopt a structured approach, decomposing decision-making into multiple layers. Here, a High-Level (HL) planner addresses navigation and path planning, while a Low-Level (LL) controller focuses on fundamental locomotion skills. The highest decision-making tier, the task level, receives direct input from task or user commands.
Figure 2: Classification of DRL-based control schemes.
@@
The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications. Although the application of RL to simple 2D bipedal robots began in 2004 [10, 11], it took several years before deep reinforcement learning (DRL) algorithms emerged. These DRL-based methods have since shown promising results in physical simulators [12, 13, 14]. Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig. 1(c) [7]. Besides model-based reference learning, the policy can also incorporate motion capture data [15, 16, 17, 18], or start from scratch [19] to explore solutions freely. Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [20, 21, 22].
@@
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. A notable framework employs a learned High-Level (HL) planner coupled with a Low-Level (LL) model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [23, 24, 25]. Another innovative construction integrates a learned feedback controller with an HL planner, categorized under the DRL Feedback Control Hybrid Scheme [26, 27]. Additionally, a learned hierarchical control scheme [28] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [12, 13, 29].
While several review papers discuss RL for general robotics [4] and model-based methods for bipeds [1, 2, 3], none specifically focus on DRL-based frameworks for bipeds. This survey aims to address this gap by summarizing current research progress, highlighting the structure and capabilities of bipedal locomotion frameworks, and exploring future directions. We also catalogue DRL-based frameworks, as depicted in Fig. 2. The primary contributions of this survey are:
@@
The paper is organized as follows: Section II focuses on end-to-end frameworks, categorized by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV addresses existing gaps, ongoing challenges, and potential future research directions. Finally, Section V concludes the paper.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly translates sensory inputs—such as images, lidar data, or proprioceptive feedback [30]—along with user commands [19] or pre-defined references [31], into joint-level control actions. These actions encompass motor torques [32], positions, and velocities [15]. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of low-level tracking to basic elements, such as a Proportional-Derivative (PD) controller. These methods can be broadly categorized based on their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
@@
Residual learning [33] [34] [35] Forward walk undirectional walk Omni-walk Add residential term to the known motor positions at the current time step. A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains. Refer to caption
@@
Reference-free learning [19] [38] [20] Periodic motions Stepping stones Vision-based Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources. Refer to caption
@@ II-A Reference-based learning
Reference-based learning utilizes prior knowledge, allowing the policy to develop locomotion skills by adhering to predefined references, which may be derived from trajectory optimization (TO) techniques or captured through motion capture systems. This method facilitates the acquisition of locomotion skills compared to alternative approaches, though it typically results in locomotion patterns that closely resemble the predefined references or motion clips, thus limiting the variety of gait patterns. Generally, this approach can be divided into two primary methods: (i) residual learning and (ii) guided learning.
@@ II-A1 Residual learning
This method involves a framework that is aware of the current reference joint positions and applies offsets determined by the policy to modify motor commands at the current timestep. By utilizing predefined motion trajectories, the residual term acts as feedback control, compensating for errors and enabling the biped to achieve dynamic locomotion skills.
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [33]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimization (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a Proportional-Derivative (PD) controller to set the desired joint positions. While this framework enhanced the robot’s ability to perform tasks beyond standing [39], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [34]. This process and its key points are further explored in Appendix B. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward. A novel approach in residual learning was introduced to enable unidirectional walking, where the policy outputs a residual term added to the current positional states, facilitating gradual omnidirectional walking [35].
@@ II-A2 Guided learning
Guided learning trains policies to directly output the desired joint-level commands, eschewing the addition of a residual term. The reward structure in this approach is focused on closely imitating predefined references.
@@
Despite these advancements, the parameterization of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [37]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviors [18]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
@@
Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities. Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
@@ II-B Reference-free learning
In reference-free learning, the policy is trained using a carefully crafted reward function rather than relying on predefined trajectories. This approach allows the policy to explore a wider range of gait patterns and adapt to unforeseen terrains, thereby enhancing innovation and flexibility within the learning process.
The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models. A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [14]. This framework introduced a novel term into the loss function and utilized a curriculum learning strategy to effectively shape gait patterns. Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [40].
Considering the practical implementation of theoretical models, significant efforts have been directed towards developing sim-to-real frameworks in robotics studies. A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [19]. This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot. It has been further refined to adapt to diverse terrains and scenarios. For instance, robust blind walking on stairs was demonstrated through terrain randomization techniques in [38]. Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [41], thus enabling the robot to effectively navigate stepping stones [20]. Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalizes terrain information [42].
@@ III Hierarchy framework
Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making. Each layer within this structure is tasked with specific objectives, ranging from high-level navigation to fundamental locomotion skills. This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
The architecture of a hierarchical framework typically comprises two principal modules: an HL planner and an LL controller. This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
@@
    Deep planning hybrid scheme: This approach combines strategic, high-level planning with dynamic low-level execution, leveraging the strengths of both learning-based and traditional model-based methods.
@@
    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with deep reinforcement learning, allowing for real-time adjustments and enhanced responsiveness.
@@
Deep Planning Hybrid Scheme [24] [43] [44] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks. A: Enhanced command tracking capabilities, generalized across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalization regarding different tasks.

### v3

# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

@@ Abstract
Bipedal robots are gaining global recognition due to their potential applications and advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL). While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge. This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes. End-to-end frameworks are evaluated based on their learning approaches, while hierarchical frameworks are examined in terms of layered structures that integrate learning-based or traditional model-based methods. We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework. Additionally, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with wide-ranging applications in real-world environments.
@@ I Introduction
Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using just two legs. Bipedal robots, which closely mimic human form, have a distinct advantage over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains. Furthermore, bipedal humanoid robots are specifically designed to function in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use. This makes them highly adaptable to a wide range of tasks in such settings.
As a result, bipedal robots hold significant potential for real-world applications [1]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [2, 3, 4]. Their agility is particularly advantageous in complex environments, such as multi-level workplaces. Bipedal robots are also well-suited for tasks that involve the use of human-designed tools, making them valuable for assisting in daily tasks, healthcare, and rehabilitation [5]. Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [9, 10, 11]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [12], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [13], which offers a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments. Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
@@
Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks? If not, how should such a framework be designed, and what features or components should it include?
To address these questions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
@@
End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers. In hierarchical systems, a High-Level (HL) planner governs navigation and path planning, while a Low-Level (LL) controller handles fundamental locomotion tasks. The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
@@
The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications. Although the application of RL to simple 2D bipedal robots began in 2004 [19, 20], it took several years before DRL algorithms emerged. These DRL-based methods have since shown promising results in physical simulators [21, 22, 23]. Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig. 1(c) [16]. Besides model-based reference learning, the policy can also incorporate motion capture data [24, 25, 26, 27], or start from scratch [28] to explore solutions freely. Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [29, 30, 31].
@@
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. A notable framework employs a learned HL planner coupled with a LL model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [32, 33, 34]. Another innovative construction integrates a learned feedback controller with an HL planner, categorised under the DRL Feedback Control Hybrid Scheme [35, 36]. Additionally, a learned hierarchical control scheme [37] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [21, 22, 38]. In order to provide a clearer understanding of the current landscape, we have categorised existing DRL frameworks, as depicted in Fig. 2.
Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure. Although bipedal locomotion tasks have evolved significantly in the past decade. Earlier frameworks primarily focused on basic walking stabilization and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation. Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [39], emphasizing real-world deployment and robustness.
Although several reviews discuss RL for general robotics [13] and model-based methods for bipedal robots [9, 10, 11], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature by applying the following selection criteria: (1) studies that examine DRL frameworks specifically for bipedal robots, (2) research involving both simulated and physical bipedal robots, (3) approaches that improve policy transfer from simulation to real-world environments, and (4) publications from the last five years (2018 - April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, ArXiv, and major robotics conferences such as CoRL, RSS, ICRA, IROS, and Humanoids.
The search was conducted using the key terms “Deep Reinforcement Learning” or “Reinforcement Learning” in combination with “Bipedal Locomotion”, “Bipedal Walking”, “Biped Robot”, “Humanoid Robot”, or “Legged Robot”. The most relevant and impactful works were manually selected for further review. This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where necessary. For clarity, in this survey, the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
The primary contributions of this survey are:
@@
The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV presents the expected unified frameworks, which build upon proposed approaches while highlighting the challenges associated with their implementation. Section V addresses potential future research directions. Finally, Section VI concludes the survey.
Figure 2: Classification of DRL-based control schemes.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, lidar data, or proprioceptive feedback [40], along with user commands [28] or pre-defined references [41], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a Proportional-Derivative (PD) controller. These methods can be broadly categorised based on their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
@@
Residual learning [42] [43] Forward walk undirectional walk Add residual term to the known motor positions at the current time step. A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains. Refer to caption
@@
Reference-free learning [28] [46] [29] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources. Refer to caption
@@ II-A Reference-based learning
Reference-based learning leverages prior knowledge generated offline through methods such as trajectory optimisation (TO) or motion capture systems. This predefined reference typically includes data related to the robot’s joint movements or pre-planned trajectories, serving as a foundation for the policy to develop locomotion skills by following these established motion patterns. Generally, this approach can be divided into two primary methods: (i) residual learning and (ii) guided learning.
@@ II-A1 Residual learning
The proposed framework utilises a policy that modifies motor commands by applying action offsets based on the current reference joint positions, allowing the biped robot to achieve dynamic locomotion through error compensation. The state space includes proprioceptive information such as trunk position, orientation, velocity, angular velocity, joint angles, and joint velocities, providing the necessary sensory data for real-time adjustments. Actions are defined by offsets, $\delta a$, which represent deviations from the predefined desired joint positions, $\hat{a}$, with the final motor commands represented as $a=\hat{a}+\delta a$. The reward function encourages the policy to optimise locomotion performance by considering (a) how closely the robot’s active joint angles match the reference angles, (b) how effectively the robot responds to user commands, and (c) additional terms that further enhance the stability of the robot’s movements. This holistic approach enables the biped robot to adapt to various dynamic conditions while maintaining balance and control.
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [42]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a PD controller to set the desired joint positions. While this framework enhanced the robot’s ability to perform tasks beyond standing [47], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [43]. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward.
A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [48]. In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive. However, this also limits the policy’s ability to explore more diverse motions, restricting it to a single slow walking pattern.
Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance. Compared to other learning approaches that directly output joint positions, it is highly sample-efficient [33]. However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
@@ II-A2 Guided learning
Guided learning trains policies to directly output the desired joint-level commands as actions $a$, without relying on the addition of a residual term. The states space is as same as residual learning approach. In this approach, the reward structure is centred on accurately imitating predefined reference trajectories, ensuring precise alignment between the policy output and the reference motion.
@@
Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [45]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [27]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
@@
Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [41, 49]. Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
@@ II-B Reference-free learning
In reference-free learning, the policy is trained using a carefully crafted reward function rather than relying on predefined trajectories. This approach allows the policy to explore a wider range of gait patterns and adapt to unforeseen terrains, thereby enhancing innovation and flexibility within the learning process. The action space and observation space in this approach are similar to the guided-learning method; however, the reward structure differs significantly from the reference-based method. Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [49].
The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models. A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [23]. This framework introduced a novel term into the loss function and utilised a curriculum learning strategy to effectively shape gait patterns. Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [50].
Considering the practical implementation of theoretical models, significant efforts have been directed towards developing sim-to-real frameworks in robotics studies. A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [28]. This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot. It has been further refined to adapt to diverse terrains and scenarios. For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [46]. Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [51], thus enabling the robot to effectively navigate stepping stones [29]. Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [52].
@@ III Hierarchy framework
Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making. Each layer within this structure is tasked with specific objectives, ranging from navigation to fundamental locomotion skills. This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
The architecture of a hierarchical framework typically comprises two principal modules: an High-Level (HL) planner and an Low-Level (LL) controller. This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
Communication between the layers in a hierarchical framework is achieved through the transmission of commands. The HL planner sets abstract goals, which the LL controller translates into specific actions, such as calculating joint movements to follow a desired trajectory. In return, the robot sends sensor data back to the HL planner, enabling real-time adjustments. The tasks handled by different layers often operate on varying time scales, adding complexity to synchronising communication between the layers.
@@
    Deep planning hybrid scheme: This approach combines strategic, HL planning with dynamic LL execution, leveraging the strengths of both learning-based and traditional model-based methods.
@@
    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with DRL, allowing for real-time adjustments and enhanced responsiveness.
@@
Deep Planning Hybrid Scheme [33] [53] [54] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks. A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.